Suppose you travel north for 35 kilometers then travel east 65 kilometers. How far are you from your starting point? North and east can be considered the directions of the y- and x-axis respectively. Round to the tenth's place.
Answer:
Show work in this space.

Answer by checkley75(3666)

LET H BE THE HYPOTENUSE. COULD HAVE USED X,Y,Z,A,B,C WHATEVER FOR THE UNKNOWN!
--------------------------------------------------------------------------------
35^2+65^2=h^2
1225+4225=h^2
5450=h^2
h=sqrt5450
h=73.8 kilometers for the distance from the starting point.